OpenStudy (anonymous):

Solve the equation the square root of the quantity x plus 4 minus 3 equals 1 for the variable. Show each step of your solution process.

OpenStudy (anonymous):



OpenStudy (freckles):

try isolating the square root part by first adding the 3 on both sides

OpenStudy (freckles):

you should then be ready to square both sides in order to cancel out the square root part

OpenStudy (marihelenh):

\[\left( \sqrt{x+4} \right)^{2}=\left( 4 \right)^{2}\]

OpenStudy (anonymous):

so the answer would be x=0?

OpenStudy (anonymous):

@freckles

OpenStudy (marihelenh):

If you were to do the previous equation, you would get \[x+4=16\]

OpenStudy (marihelenh):

Then just subtract across and get x and finish with a check.

OpenStudy (anonymous):

oh yes i forgot to square it thanks

OpenStudy (anonymous):

so x=12

OpenStudy (marihelenh):

Yep! And it checks.

OpenStudy (anonymous):

thank you @marihelenh